OpenStudy (anonymous):

What is the geometric mean of: 3 x - = - x 32

myininaya (myininaya):

I'm afraid I don't get the question

OpenStudy (anonymous):

It says "solve for", but the section is referring to the geometric mean.

OpenStudy (mertsj):

x is the geometric mean of 3 and 32 according to your setup.

myininaya (myininaya):

so we are suppose to solve for ok great cross multiply

myininaya (myininaya):

\[3(32)=x^2\]

OpenStudy (anonymous):

but 64 isnt the correct answer, the choices are " 96 48 96 48

myininaya (myininaya):

now take the square root of both sides

myininaya (myininaya):

\[x^2=96\]

myininaya (myininaya):

\[x=\sqrt{96}\]

OpenStudy (anonymous):

the choices are A) 96 B) 48 C) 6 Square root 4 D) 4 Square root 6 so the answer would be 96?

myininaya (myininaya):

there are perfect squares in 96

OpenStudy (mertsj):

No. 96 = 16x6 so 4sqrt6

OpenStudy (anonymous):

Ohhhhh. Tysm! I get it more since you explained it.The book sucks at explaining it.

myininaya (myininaya):

\[\sqrt{96}=\sqrt{16 \cdot 6}=\sqrt{16} \sqrt{6}\]